Sue had 6 times as many quarters as Jim. Ben has twice as many quarters as Jim. Together they had $9. How much money did each person have?
Answer by ewatrrr(24785)


Hi
Let x, 6x, 2x represent Jim's, Sues's and Ben's quarters respectively
Question states***CENTS makes sense
25*9x = 900
Solving for x
x = 4, number of quarters Jim had. Sue 24 and Ben 8
CHECKING our Answer***
25*(4+24+8) = 25*36 = 900 or $9.00